# Implementing LLMs in Industrial Process modeling: Addressing categorical variables

### v1

@@ Abstract
Important variables of processes are, in many occasions, categorical, i.e. names or labels representing, e.g. categories of inputs, or types of reactors or a sequence of steps. In this work, we use Large Language Models (LLMs) to derive embeddings of such inputs that represent their actual meaning, or reflect the “distances" between categories, i.e. how similar or dissimilar they are. This is a marked difference from the current standard practice of using binary, or one-hot encoding to replace categorical variables with sequences of ones and zeros. Combined with dimensionality reduction techniques, either linear such as Principal Components Analysis (PCA), or nonlinear such as Uniform Manifold Approximation and Projection (UMAP), the proposed approach leads to a meaningful, low-dimensional feature space. The significance of obtaining meaningful embeddings is illustrated in the context of an industrial coating process for cutting tools that includes both numerical and categorical inputs. The proposed approach enables feature importance which is a marked improvement compared to the current state-of-the-art (SotA) in the encoding of categorical variables.
@@ 1 Introduction
When dealing with industrial/production data, it is often the case that data are not numerical values, e.g. temperature time-series, flow-rates, pressure etc., but rather involve categorical variables, i.e. categories described by an assigned name. Examples include serial numbers of reactors, working-names of products or even an entire sequence of steps represented by a single name. One of the most popular methods to address this issue is one-hot encoding, which replaces categorical variables with several columns of dummy variables, commonly a sequence of zeros and ones. By replacing categorical variables with numerical ones, it is possible to use them as features in various machine learning algorithms. However, by doing so, the resulting representation is agnostic of the actual meaning of the variables, and it is unable to reflect the actual "distance" (i.e. relative similarity or dissimilarity) between the different categories. For these reasons, it is an obstacle for feature importance and sensitivity analysis, uncertainty quantification, and/or optimization, diminishing the opportunity to create insights from the data and increase understanding of the physical process at hand.
Following their rise to popularity, a lot of interest has fallen in LLMs and their implementation [11, 12]. Specifically in chemical and process engineering, the prospects of LLMs for education [13] and process mining [14] have been examined. With their work, Decardi-Nelson et al. [15] investigate the implementation of the generative ability of LLMs in the various stages of process systems engineering (PSE), namely synthesis and design, optimization and integration, and process monitoring and control. In [16], the use of an agent-based LLM approach in the context of process simulation is proposed, combining four different LLM-based agents for the observation of the process. Furthermore [17] evaluate the capabilities of LLMs for the prediction, generation and understanding of tabular data. However, despite the widespread use of LLMs for text embedding, exemplified by S-Bert [18], to the best of our knowledge, the impact of using LLMs to encode categorical variables in the development of predictive models has not been explored.
@@
With this study, we aim to explore the potential benefits of implementing LLMs to derive meaningful embeddings of categorical data, using as inputs short textual descriptions of what the different categories represent. These embeddings are then compressed using either linear or nonlinear dimensionality reduction methods and used as inputs to tree-based regression models. Actual production data are used in this study, stemming from an industrial chemical vapor deposition reactor used for the production of wear-resisting coatings on cutting tools, called “inserts". This process combines competing physical and chemical mechanisms in very complex and ever-changing geometrical set-ups. There are no inline measurements and therefore there is no possibility to monitor the process as it happens and intervene with control actions. Coating thickness measurements are collected ex-situ, after the process is finished, providing information about the quality of the product.
@@
Aiming to develop a data-driven predictive model [19] for the quality characteristics of the coating, a tree-based regression algorithm is trained using several numerical and categorical inputs. The categorical inputs include the insert (cutting tool) name, a series of letters and numbers, representing an object with an ISO-specified shape and size. In what we call the “original" (SotA) treatment of the data, the entire name is encoded with one-hot encoding, which is oblivious of the meaning of each component of the name. This poses a restriction on the amount of information that may be extracted from the data-driven model. As an example, it is not possible to establish if and what effect the distribution of inserts in the reactors, their placement, in general, has on the output, i.e. the quality of the product.
@@
In this work, several models, pre-trained an un-trained, are compared and subsequent feature importance analysis, using Shapley values, reveals that critical process inputs; physically meaningful and useful conclusions arise when using the proposed embedings in contrast to binary encoding.
@@
The development of the proposed tool leads to improved understanding of the process set up since it paves the way for sensitivity analysis and uncertainty quantification. Furthermore, it allows the incorporation into datasets of pieces of production information that is maintained in the form of notes, short texts, in the production diaries. This information is often very important and is assessed by experts in the production line but not in a systematic way.
@@ 2 Process description
The primary objective of the process is to achieve a uniform coating thickness, as this uniformity directly correlates with consistent product longevity [23]. Ideally, this uniformity would be consistent across all production runs, reactors, and sites. However, this is not always achieved. Therefore, it is crucial to develop a method for predicting the coating thickness of the inserts based on the reactor setup. In addition, establishing a systematic approach to evaluate factors affecting the uniformity of the coating thickness is essential. To this extent, the application of both equation-based methods [20] and data-driven methods [21, 19] has been demonstrated in previous work, to which the interested reader is referred for further information on the process.
@@ 3 The original data-driven approach
In addition to coating thickness measurements, the dataset also contains several features regarding the set-up of the production run. These features are used as inputs for our predictive ML models. An important feature is the production “recipe”, which encapsulates the steps taken and the process conditions during production. Furthermore, there are features that contain information regarding the “how” and the “where” the inserts are placed within the reactor. More specifically, in the original approach, the following features are used: a) The number of inserts placed on each disk. b) The position of each disk within the reactor. c) The type of insert placed on each disk. Each type of insert has different geometrical characteristics. d) The surface area of the inserts placed on each disk.
@@ 4 Methods
This section provides a succinct description of the methods and algorithms used to create vector representations of insert shape descriptions through various embeddings, and to present similarities between the different representations. These vectors are included as numerical features in a regression model that predicts insert thickness based on other insert features listed in Section 3. Additionally, the model performance will be evaluated and compared across different embeddings using a cross-validation method, followed by feature importance and Shapley analysis.
@@ 4.1.1 Doc2Vec
Doc2Vec [31] [32] is an unsupervised machine learning algorithm from the Gensim library to obtain vector representations of documents. In this case, a document is an object of the text sequence type and it could be anything from a short character string to documents structured by paragraphs such as an article or a book.
This algorithm is an extension of the Word2Vec, based on the bag-of-words [28] and bag-of-n-grams [28, 33]. It was designed to create fixed-length vector representations from pieces of texts. This model allows us to capture semantic meanings (distances between the words) and context (relationships between words and documents), which increase the accuracy of various tasks such as document classification, clustering, and recommendation.
@@ 4.1.4 Similarity between descriptions
Cosine similarity is a widely used metric in NLP for assessing the similarity between two vectors. In the context of word embeddings, cosine similarity is a valuable method for comparing the similarity of words or documents [39, 40].
It measures the cosine of the angle between two vectors in a high-dimensional space. The similarity score ranges from -1 to 1, where a score of -1 signifies completely opposite vectors, 0 indicates no similarity, and 1 means identical vectors. The cosine similarity between two vectors is computed using the following formula:
where represents the dot product between vectors and , while denotes the Euclidean norm.
@@ 4.2.1 K-fold cross-validation
Therefore, this approach is more efficient than using a fixed validation set, as it reduces computational cost and makes better use of the available data.
@@ 5.1 Data preprocessing
Regarding the insert names, each ISO designation code (cf. Fig. 3) is broken down into its various features, aiming to provide additional information that could be useful to include in the XGBOOST model.
@@
The new features are summarized in Table 3). In the data transformation stage, we first apply data imputation to ensure that missing values were not included in our analysis. Then, we standardize the numerical features and use embeddings/transformers to represent categorical features as vectors.
@@ 5.2 Similarity
Concretely, for the Doc2Vec model (see Fig. 4), we first observe a sort of reflexive property between descriptions, meaning each description is perfectly similar to itself. Additionally, we can see certain groups of descriptions that are expected to be similar, such as the group of rhombus-shaped inserts with different degrees. We also notice specific connections between certain shapes, like triangular and circular shapes with the rounded triangular shape, or some similarity among standard shapes like triangles, circles, and squares. It is also interesting to note that the filling material is close to the unknown inserts shape. This illustrates that the Doc2Vec model is capturing reasonably well the semantic and contextual meaning of the descriptions.
Additionally, in the case of the all-MiniLM-L12-v2 model, we again observe the reflexivity property. We also see similar patterns to those of the previous case, but more marked (refer to Fig. 5), with values surpassing 0.83, among semantically close insert shape descriptions, as expected. For instance, cosine_similarity values for rhombus or triangular shapes. Two additional observations stand out: the corresponding similarity values for the descriptions filling material and unknown shape with respect to other descriptions are below 0.44 in the first case and closer to zero in the second case. This indicates that the model is correctly capturing their meanings, and we have a reasonably accurate representation of the considered descriptions.
For the all-mpnet-base-v2 model, we obtain results similar to those of the all-MiniLM-L12-v2 model, maintaining the same patterns (see Fig. 6). Once again, the values for descriptions such as filling material or unknown shape are distinct from other groups, with values below 0.41. Additionally, there is a high similarity among the descriptions in the group of quadrilaterals located in the upper left of Fig. 6, with values exceeding 0.88. However, there is no clear differentiation between standard shapes such as squares, circles, or triangles, which have values in the middle and lower right part of this Fig. 6.
@@
In contrast, the second scenario corresponds to the high dimensionality of the latent spaces produced by both transformers, possible consequences of reducing their dimensionality can be explored. For this, we consider a linear method PCA of dimension 3. The results obtained are presented in Appendix A and show the same patterns observed when dimensionality reduction is not applied and capturing the same semantic and contextual meaning, as shown in Figs. 13 and 14. In the next section, we will also analyze how both cases impact the model’s performance.
@@ 5.3 XGBOOST model performance
In all cases, the vectors obtained by the transformers are included as numerical features in the XGBOOST model. At this point, it is interesting to explore how the amount of information considered during training influences the model’s performance. To address this, we will examine two important metrics in supervised machine learning models that provide insight into their accuracy and robustness: mean squared error (MSE) and score.
@@ 5.4 Mean squared error, MSE
In the first case, we compare the original model, which shows a decreasing linear trend as the training data size increases (black dashed line). However, its confidence interval also narrows after considering 80% of the training data, suggesting reduced variability around the MSE mean and greater confidence that the true MSE value lies within this (black) interval. Comparing all intervals, this would indicate possible limitations in further reducing the error for the original model. When examining the curves for the cases where Doc2Vec is used to encode either only the insert shapes (in blue) or both the insert shapes and the recipe (in green), we observe that the average MSE follows a similar decreasing trend. Additionally, we see a narrowing of their respective confidence intervals after considering 70% of the data, and these intervals cover a slightly lower range than the original model’s, suggesting that further reduction in error may be possible, as the bounds of both confidence intervals (blue and green) are slightly below that of the original model. This could also be achieved by incorporating more data or by fine-tuning the model’s parameters.
In the other two cases, corresponding to the application of all-mpnet-base-v2 and all-MiniLM-L12-v2 transformers, we observe similar behavior to Doc2Vec, except that the average error in each case is very close to the average MSE of the original model, indicating no significant improvements. Furthermore, in the case of transformers, considering 80% of the training data results in a widening of their confidence intervals, suggesting that these representations require as much data as possible to achieve performance equal to or only marginally better than that of the original model. This approach could reduce the variability of the MSE mean, resulting in narrower intervals and providing greater confidence that the true value falls within this range.
5.5 score
Similarly to the previous section, we will analyze the evolution of the metric as a function of the proportion of data used for training, through -fold cross-validation.
To assess whether the obtained values follow a normal distribution, we again use the Kolmogorov-Smirnov test. As shown in Table 5, the p-values (> 0.05) support the assumption of normality.

### v2

# Implementing NLPs in Industrial Process modeling: Addressing categorical variables

@@ Abstract
Important variables of processes are often categorical, i.e. names or labels representing, e.g. categories of inputs, or types of reactors or a sequence of steps. In this work, we use Natural Language Processing Models to derive embeddings of such inputs that represent their actual meaning, or reflect the “distances" between categories, i.e. how similar or dissimilar they are. This is a marked difference from the current standard practice of using binary, or one-hot encoding to replace categorical variables with sequences of ones and zeros. Combined with dimensionality reduction techniques, either linear such as Principal Component Analysis, or nonlinear such as Uniform Manifold Approximation and Projection, the proposed approach leads to a meaningful, low-dimensional feature space. The significance of obtaining meaningful embeddings is illustrated in the context of an industrial coating process for cutting tools that includes both numerical and categorical inputs. In this industrial process, subject matter expertise suggests that the categorical inputs are critical for determining the final outcome but this cannot be taken into account with the current state-of-the-art. The proposed approach enables feature importance which is a marked improvement compared to the current state-of-the-art in the encoding of categorical variables. The proposed approach is not limited to the case-study presented here and is suitable for applications with similar mix of categorical and numerical critical inputs.
@@ 1 Introduction
When dealing with industrial/production data, it is often the case that data are not numerical values, e.g. temperature time-series, flow-rates, pressure etc., but rather involve categorical variables, i.e. categories described by an assigned name. Examples include serial numbers of reactors, working-names of products or even an entire sequence of steps represented by a single name. One of the most popular methods to address this issue is one-hot encoding [11, 12, 13], which replaces categorical variables with several columns of dummy variables, commonly a sequence of zeros and ones. By replacing categorical variables with numerical ones, it is possible to use them as features in various machine learning algorithms. However, by doing so, the resulting representation is agnostic of the actual meaning of the variables, and it is unable to reflect the actual "distance" (i.e. relative similarity or dissimilarity) between the different categories. For these reasons, it is an obstacle for feature importance and sensitivity analysis, uncertainty quantification, and/or optimization, diminishing the opportunity to create insights from the data and increase understanding of the physical process at hand.
Following their rise to popularity, a lot of interest has fallen in Natural Language Processing (NLP) models and their implementation [14, 15], because these models can generate embeddings that represent categorical variables as dense vectors of real numbers, which can be incorporated into other computational tasks. Specifically in chemical and process engineering, the prospects of NLP, for materials characterization [16] and text mining [17, 18, 19] have been examined. In [20], the use of an NLP approach in the context of vapor deposition process is proposed, combining a systems-based approach with NLP to distinguish essential mechanisms and efficiently extract causal knowledge from extensive literature. Furthermore, [21] systematically examines the development of language modeling for tabular data, covering data structures and types, key datasets and evaluation tasks, modeling techniques and the evolution of pre-trained models. However, despite the widespread use of NLP for text embedding, exemplified by S-Bert [22], to the best of our knowledge, the impact of using Language models to encode categorical variables in the development of predictive models has not been explored.
@@
With this study, we aim to explore the potential benefits of implementing NLPs to derive meaningful embeddings of categorical data, using as inputs short textual descriptions of what the different categories represent. These embeddings are then compressed using either linear or nonlinear dimensionality reduction methods and used as inputs to tree-based regression models. Actual production data are used in this study, stemming from an industrial chemical vapor deposition reactor used for the production of wear-resisting coatings on cutting tools, called “inserts". This process combines competing physical and chemical mechanisms in very complex and ever-changing geometrical set-ups. There are no inline measurements and therefore there is no possibility to monitor the process as it happens and intervene with control actions. Coating thickness measurements are collected ex-situ, after the process is finished, providing information about the quality of the product.
@@
Aiming to develop a data-driven predictive model [23] for the quality characteristics of the coating, a tree-based regression algorithm is trained using several numerical and categorical inputs. The categorical inputs include the insert (cutting tool) name, a series of letters and numbers, representing an object with an ISO-specified shape and size. In what we call the “original" state-of-the-art (SotA) treatment of the data, the entire name is encoded with one-hot encoding, which is oblivious of the meaning of each component of the name. This poses a restriction on the amount of information that may be extracted from the data-driven model. As an example, it is not possible to establish if and what effect the distribution of inserts in the reactors, their placement, in general, has on the output, i.e. the quality of the product.
@@
In this work, several models, pre-trained an un-trained, are compared and subsequent feature importance analysis, using Shapley values, reveals that critical process inputs; physically meaningful and useful conclusions arise when using the proposed embeddings in contrast to binary encoding.
@@
The development of the proposed tool leads to improved understanding of the process set up since it paves the way for sensitivity analysis and uncertainty quantification. Furthermore, it allows the incorporation into datasets of pieces of production information that are maintained in the form of notes, short texts, in the production diaries. This information is often very important and is assessed by experts in the production line but not in a systematic way.
@@ 2 Process description
The primary objective of the process is to achieve a uniform coating thickness, as this uniformity directly correlates with consistent product longevity [27, 28]. Ideally, this uniformity would be consistent across all production runs, reactors, and sites. However, this is not always achieved. Therefore, it is crucial to develop a method for predicting the coating thickness of the inserts based on the reactor setup. In addition, establishing a systematic approach to evaluate factors affecting the uniformity of the coating thickness is essential. To this extent, the application of both equation-based methods [24] and data-driven methods [25, 23] has been demonstrated in previous work, to which the interested reader is referred for further information on the process.
@@ 3 The original data-driven approach
In addition to coating thickness measurements, the dataset also contains several features regarding the set-up of the production run. These features are used as inputs for our predictive ML models. An important feature is the production “recipe”, which encapsulates the steps taken and the process conditions during production. These specific details cannot be detailed here. Furthermore, there are features that contain information regarding the “how” and the “where” the inserts are placed within the reactor. More specifically, in the original approach, the following features are used: a) The number of inserts placed on each disk. b) The position of each disk within the reactor. c) The type of insert placed on each disk. Each type of insert has different geometrical characteristics. d) The surface area of the inserts placed on each disk.
@@
The XGBOOST model (cf. Section 4.3) is the most efficient and high-performing ML model for this task, as implemented and shown in [23]. Therefore, this model will also be considered for the present work.
@@ 4 Methods
This section provides a succinct description of the methods and algorithms used to create vector representations of insert shape descriptions through various embeddings, and to present similarities between the different representations. These vectors are included as numerical features in a predictive model that predicts insert thickness based on other insert features listed in Section 3. Additionally, the model performance will be evaluated and compared across different embeddings using a cross-validation method, followed by feature importance and Shapley analysis.
@@ 4.1 Embeddings and transformers
Existing studies evaluated these models on various benchmark problems [36, 37, 38, 6, 39], consistently demonstrating strong performance, especially in sentence similarity tasks. The results and selected pre-trained models are available in the Hugging Face open-access transformers repository, where the top two models were chosen based on performance scores [40]
Non-pretrained models like Doc2Vec [41] generate similar vector representations for related categories, enhancing categorical variable integration in machine learning, especially in industrial settings where these variables reflect key interdependencies [42] (e.g., material types, operating modes, tool geometries). Doc2Vec also allows control over latent space dimensionality, balancing efficiency and performance. Meanwhile, pre-trained models offer robust sentence embeddings that capture contextual relationships. Chosen for their embedding quality, efficiency, and scalability, these models effectively categorize and cluster industrial categorical variables [43, 44].
@@ 4.1.1 Doc2Vec
Doc2Vec [45] is an unsupervised machine learning algorithm from the Gensim library to obtain vector representations of documents. In this case, a document is an object of the text sequence type and it could be anything from a short character string to documents structured by paragraphs such as an article or a book.
This algorithm is an extension of Word2Vec, based on the bag-of-words [33] and bag-of-n-grams [33, 46]. It was designed to create fixed-length vector representations from pieces of texts. This model allows us to capture semantic meanings (distances between the words) and context (relationships between words and documents), which increase the accuracy of various tasks such as document classification, clustering, and recommendation.
@@ 4.1.4 Similarity between descriptions
Cosine similarity is a widely used metric in NLP for assessing the similarity between two vectors. In the context of sentences embeddings, cosine similarity is a valuable method for comparing the similarity of words or documents [47, 12].
It measures the cosine of the angle between two vectors in a high-dimensional space. The similarity score ranges from -1 to 1, where a score of -1 signifies completely opposite vectors, 0 indicates no similarity, and 1 means identical vectors. The cosine similarity (Sim) between two vectors is computed using the following formula:
where represents the dot product between vectors and , while denotes the Euclidean norm. In Fig. 4, the calculation of cosine similarity is illustrated, along with the pairwise computation of the similarity scores derived from the vector representations generated by embeddings and transformers.
4.2 Dimensionality reduction
To effectively handle the high dimensionality of embedding spaces generated by different transformers (cf. Sections 4.1 and 5), we employ linear dimensionality reduction methods, such as Principal Component Analysis (PCA) [48, 49], which is based on singular value decomposition. This technique reduces the original data while preserving its overall variability. Such an approach has been successfully applied for dimensionality reduction in the context of CVD processes [50, 51]. Alternatively, nonlinear methods like Uniform Manifold Approximation and Projection (UMAP) [52, 53], a manifold learning technique for dimensionality reduction, can be employed. UMAP is also widely used in chemical engineering for data visualization [54, 55].
@@ 4.3.1 K-fold cross-validation
Therefore, this approach is more efficient in treating the available data and addressing the randomness of the test-set selection.
@@ 5.1 Data preprocessing
Regarding the insert names, each ISO designation code (cf. Fig. 3) is broken down into its various features, aiming to provide additional information that could be useful to include in the XGBOOST model. Additionally, it is important to note that the other categorical feature, production “recipe”, is independent of the inset shape. This is because the reactor contains various inserts with different shapes, making this feature irrelevant when employing embeddings.
@@
The new features are summarized in Table 3. In the data transformation stage, we first apply data imputation to ensure that missing values were not included in our analysis. Then, we standardize the numerical features and use embeddings/transformers to represent categorical features as vectors.
@@ 5.2 Similarity
As an example, let us consider the following cases. In Fig. 5, the value enclosed in the yellow square indicates that the fifty-five-degree rhombus has a high similarity in shape characteristics to the thirty-five-degree rhombus. In the case of intermediate values, such as the one enclosed in the green square, it signifies that the thirty-five-degree rhombus exhibits some common shape attributes with the rectangular insert. Finally, the lower value, enclosed in a blue square, confirms a significantly different shape description, illustrating the distinction between the unknown insert shape and an eighty-five-degree rhombus insert shape.
Concretely, for the Doc2Vec model (see Fig. 5), we first observe a sort of reflexive property between descriptions, meaning each description is perfectly similar to itself. Additionally, we can see certain groups of descriptions that are expected to be similar, such as the group of rhombus-shaped inserts with different degrees. We also notice specific connections between certain shapes, like triangular and circular shapes with the rounded triangular shape, or some similarity among standard shapes like triangles, circles, and squares. It is also interesting to note that the filling material is close to the unknown inserts shape. In this context, the filling material refers to a non-specified insert shape that occupies the void in a disc, ensuring uniformity during the deposition process and maintaining the structural integrity of the disc. Therefore, these results illustrate that the Doc2Vec model is capturing reasonably well the semantic and contextual meaning of the descriptions.
Additionally, in the case of the all-MiniLM-L12-v2 model, we again observe the reflexivity property. We also see similar patterns to those of the previous case, but more marked (refer to Fig. 6). In the with values surpassing 0.83, among semantically close insert shape descriptions, as expected. For instance, in Fig. 6 (figure above) cosine_similarity values for rhombus or triangular shapes. Two additional observations stand out: the corresponding similarity values for the descriptions filling material and unknown shape with respect to other descriptions are below 0.44 in the first case and closer to zero in the second case. This indicates that the model is correctly capturing their meanings, and we have a reasonably accurate representation of the considered descriptions.
For the all-mpnet-base-v2 model, we obtain results similar to those of the all-MiniLM-L12-v2 model, maintaining the same patterns (see Fig. 6). In Fig. 6 (figure bellow), the values for descriptions such as filling material or unknown shape are distinct from other groups, with values below 0.41. Additionally, there is a high similarity among the descriptions in the group of quadrilaterals located in the upper left of Fig. 6, with values exceeding 0.88. However, there is no clear differentiation between standard shapes such as squares, circles, or triangles, which have values in the middle and lower right part of this Fig. 6.
@@
In contrast, the second scenario corresponds to the high dimensionality of the latent spaces produced by both transformers. Possible consequences of reducing their dimensionality can be explored. For this study, we apply a linear dimensionality reduction method, PCA, to compress the data into three dimensions (based on an analysis of the reconstruction error), and a non-linear method, UMAP, with the same dimensionality. Although UMAP is typically more efficient for dimensionality reduction, in this case, PCA provides a better representation by preserving the semantic and contextual understanding captured by the embeddings. This is illustrated in Fig. 7 for both transformers: all-MiniLM-L12-v2 and all-mpnet-base-v2, where similar patterns are observed compared to when no dimensionality reduction is implemented. The results for UMAP are presented in Appendix A.1, where we observe slightly different patterns, suggesting that UMAP may still be a better fit for other datasets.
In the next section, we will also analyze how both scenarios impact the predictive model’s performance.
@@ 5.3 XGBOOST model performance
In all cases, the vectors obtained by the transformers are included as numerical features in the XGBOOST model, as this algorithm demonstrated the best performance in the original implementation presented in [23]. At this point, it is interesting to explore how the amount of information considered during training influences the model’s performance. To address this, we will examine two important metrics in supervised machine learning models that provide insight into their accuracy and robustness: mean squared error (MSE) and score.
@@ 5.4 Mean squared error, MSE
In Fig. 8 (a), we illustrate the results for the original model, which shows a decreasing linear trend as the training data size increases (black dashed line). However, its confidence interval also narrows after considering 80% of the training data, suggesting reduced variability around the MSE mean and greater confidence that the true MSE value lies within this (black) interval. Comparing all intervals, this would indicate possible limitations in further reducing the error for the original model. When examining the curves for the cases where Doc2Vec is used to encode either only the Insert shape features (in blue) or both the Insert shape features and the recipe (in green), we observe that the average MSE follows a similar decreasing trend. Additionally, we see a narrowing of their respective confidence intervals after considering 70% of the data, and these intervals cover a slightly lower range than the original model’s, suggesting that further reduction in error may be possible, as the bounds of both confidence intervals (blue and green) are slightly below that of the original model. This could also be achieved by incorporating more data or by fine-tuning the model’s parameters.
In the other two cases, corresponding to the application of all-MiniLM-L12-v2 (Fig. 8 (b) and (b’)) and all-mpnet-base-v2 (Fig. 8 (c) and (c’)) transformers with and without the implementation of PCA, respectively), we observe similar behavior to Doc2Vec, except that the average error in each case is very close to the average MSE of the original model, indicating no significant improvements. Furthermore, in the case of transformers, considering 80% of the training data results in a widening of their confidence intervals, suggesting that these representations require as much data as possible to achieve performance equal to or only marginally better than that of the original model. This approach could reduce the variability of the MSE mean, resulting in narrower intervals and providing greater confidence that the true value falls within this range.
5.5 R2 score
Similarly to the previous section, we will analyze the evolution of the metric as a function of the proportion of data used for training, through -fold cross-validation. Then, to assess whether the obtained values follow a normal distribution, we again use the Kolmogorov-Smirnov test. As shown in Table 5, the p-values (> 0.05) support the assumption of normality.